Sodium azide is known to inhibit Cytochrome c Oxidase. Using sodium azide at a concentration of 2 µM, you run an enzyme assay. After careful data analysis, you determine the following values.

Vmax = 50 µM/sec                                            Vmax(app) = 50 µM/sec

Km = 6 µM                                                      Km (app) = 18 µM

  1. Determine the type of inhibition present.
  1. Calculate the associated Ki(a), and (or) Ki(b), if applicable. Show your work.
Answer #1

A- Given that Vmax remains, the same but Km is increased. In the case of competitive inhibition, Vmax remains the same and Km increase. In the case of non-competitive inhibition, Km remains the same but Vmax decreases. In the case of uncompetitive inhibition both changes.

So the type of inhibition is competitive inhibition.

B- Kmapp = Km (1+[I]/Ki)

Here Kmapp = 18uM

Km = 6uM

I = 2uM

Ki need to find out

18 = 6(1+2/Ki)

18/6 = (Ki+2)/Ki

3Ki = Ki+2

3Ki-Ki =2

Ki = 2uM

So Ki = 2uM